A flask is charged with 2.00 atm of nitrogen dioxide and 1.00 atm of dinitrogen tetroxide at 25 °C and allowed to reach equilibrium. When equilibrium is established, the partial pressure of NO2 has decreased by 1.24 atm. 

(a) What are the partial pressures of NO2 and N2O4 at equilibrium? 

(b) Calculate and for Kp and Kc following reaction at 25 °C.

2NO2 (g) ↔ N2O2 (g)

(Answer: K = 2.80; K c = 68.6)


1
Expert's answer
2022-04-19T03:11:59-0400


 


Kp = 0.38 / (0.76)2 = 0.658

Kc = Kp / (RT)Δn

Δn = 1 – 2 = -1

T = 25C = 298 K

Kc = 0.658 / (8.314 * 298)-1 = 1630